Part 1:  Make or Buy Decisions

Mobility Partners makes wheelchairs and other assistive devices. For years it has made the rear wheel assembly for its wheelchairs. A local bicycle manufacturing firm, Trailblazers, Inc., offered to sell these rear wheel assemblies to Mobility. If Mobility makes the assembly, its cost per rear wheel assembly is as follows (based on annual production of 2,000 units):

                                     

Direct materials          $          30         

Direct labor                             52         

Variable overhead                   21         

Fixed overhead                       52         

Total                             $           155     

Trailblazers has offered to sell the assembly to Mobility for $110 each. The total order would amount to 2,000 rear wheel assemblies per year, which Mobility's management will buy instead of make if Mobility can save at least $20,000 per year. Accepting Trailblazers's offer would eliminate annual fixed overhead of $41,100.

Required:

a. Using Excel, prepare a schedule that shows the total differential costs. Indicate if the differential cost is higher, lower or no effect under the assumption that Status Quo is the base.  

b. Should Mobility make rear wheel assemblies or buy them from Trailblazers?  Why? Be sure to provide a complete justification for your decision.
Answer #1

Answer:

a.

Status Quo Alternative Difference
Trailblazers’ offer          220,000          220,000 higher
Materials            60,000            60,000 lower
Labor          104,000          104,000 lower
Variable Overhead            42,000            42,000 lower
Fixed Overhead applied          104,000            62,900            41,100 lower
Total Costs          310,000          282,900          (27,100) lower

Calculation of above values:

Status Quo = Based on 2,000 units

Fixed overhead applied = $52 × 2,000 units = $104,000; 104,000− $41,100 = $62,900

Trailblazers’ offer Alternative = 110*2000 = 220,000

Status Quo Costs incurred :

Labor = 52 * 2000 = 104,000

Materials = 30*2000 = 60000

Variable overhead = 21*2000 = 42000

Fixed overhead = 52*2000 = 104,000

b. Buy

Mobility could buy rear wheel assembles from Trailblazers, than making those. The $20,000 savings could be achieved by buying that. The cost to make is $27,100 more than the cost to purchase from Trailblazers.